Assume the annual retention rate for a cell phone subscriber is 70 percent and the customer generates $300 per year in profit. Assuming an annual discount rate of 8 percent, compute the value of a customer.
Answer #1

The value of the customer can be calculated by considering the profit they generate, retention rate, and the discount.

Value = Profit per year * retention rate * (1 - discount)

Value = 300 * 0.7 * (1 - 0.08) = 300 * 0.7 * 0.92 = 193.2

The value of a customer is $193.2.